What is the final pressure inside a scuba tank after it cools from 25°C to -10°C? The initial pressure in the

tank is 130 atm. 


1
Expert's answer
2022-04-22T08:08:04-0400

T1 = 250C = 298 K

T2 = -100C = 263 K

P1 / T1 = P2 / T2

130 / 298 = P2 / 263

P2 = 114.7 atm